OpenStudy (anonymous):

QUESTION 5 OF 10 The numbered disks shown are placed in a box and one disk is selected at random. Find the probability of selecting a green disk, given that an even number is selected. Write your answer as a fraction in simplest form.

OpenStudy (anonymous):

... how do you impose i do this.

OpenStudy (anonymous):

Depends on how many disks and what colors?

OpenStudy (anonymous):

There are eight disks 4 are green and 4 are blue

OpenStudy (anonymous):

Green are 3,4 ,5,7

OpenStudy (anonymous):

Okay so the probability of selecting a green disk is: \[\frac{ 4 }{ 8 }\] which, simplified, is: \[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Because, Since there are 4 green disks, and 4 blue disks, then all together there are 8 disks. And the probability of selecting a green disk depends on how many green disks there are. There are 4 so, \[\frac{ 4 }{ 8 }\] which simplifies into one half.

OpenStudy (anonymous):

It wasnt right

OpenStudy (anonymous):

What did you mean by "Green are 3, 4, 5, 7"?

OpenStudy (anonymous):

Read the initial question i think the green disk is supposed to be even